OpenStudy (anonymous):

Solve -2x + 1 -7 and describe the graph of the solution Please HELP

OpenStudy (anonymous):

-2x+1-7 2x+6 x=-3 no graph bcs your eqn ?

OpenStudy (anonymous):

i assume question is -2x+1-7=0 add seven on both sides -2x+1-7+7=0+7 -2x+1=7 subtract 1 from both sides -2x-1+1=7-1 -2x=6 divide both sides by -2 x=-3

OpenStudy (anonymous):

Choices were; 1) x 4; closed circle on 4, shading to the left 2) x 4; closed circle on 4, shading to the right 3) x 4; open circle on 4, shading to the right 4) x 4; open circle on 4, shading to the left